# Multi-Task Models Adversarial Attacks

### v3

@@ Appendix B MTL Attack Optimization Approximation
As the problem is intractable, we make some approximations and reformulate it be an Integer Linear Programming (ILP) problem. To do so, we first apply the Taylor expansion on in the numerator of the objective function at the point of :
@@
Appendix C Algorithm Pseudocode
As presented in Section 3 in the main paper, integrating GB-MTA with any existing attack can be easily accomplished by substituting the single-task gradient, i.e., , with the balanced multi-task counterpart, i.e., . We illustrate how to integrate GB-MTA into a single-task attack algorithm APGD and a multi-model attack algorithm Auto-SAGE.
APGD and GB-MTA-APGD
We provide pseudocode comparisons for APGD with and without integrating GB-MTA in Algorithms C and C. We color the difference in the two algorithms in blue. To integrate GB-MTA in APGD, we first change the key part of the code determining the attack direction (lines 2 and 11) to the proposed balanced multi-task gradients. Then we further update the objective function (lines 4 and 14) from the absolute loss value to the relative loss value sum over all tasks to accommodate the multi-task setting. All other lines in Algorithm C are kept the same as the original APGD.
Algorithm 2 APGD
Algorithm 3 GB-MTA-APGD
Auto-SAGE and GB-MTA-Auto-SAGE
For Auto-SAGE (Rathbun et al. 2022), the original attack formulation is
To integrate GB-MTA with Auto-SAGE, we normalize the gradient of the objective function and with the objective function value. The updated attack will be,
Appendix D More Experimental Results
This section reports evaluation metrics and more experimental results that are ommited from the main paper.
Evaluation Metrics
@@ Appendix D More Experimental Results
Results of Attack Performance
@@
Figure 5 illustrates the attack performance on NYUv2 with MobileNetV2. To be consistent with the main paper, we conduct the experiments using different variants of GB-MTA and the naïve multi-task attacks. The x-axis represents the perturbation bound ranging from 1 to 16, while the y-axis displays the overall Average Relative Performance (ARP, higher-the-better). Overall, GB-MTA consistently outperforms baselines.
@@
We present the full tables of ARP after the attack of all 25 multi-task models trained on NYUv2 with Deeplab-ResNet34 for perturbation bound in Tables 4 and 5. Similarly, Tables 6 and 7 show the attack results for all 15 multi-task models for Tiny-Taskonomy. Overall, GB-MTA achieves 80% first place (80 out of 100 cases) on NYUv2 and 88.33% (53 out of 60) on Tiny-Taskonomy, demonstrating the effectiveness of adversarial samples from GB-MTA.

### v4

@@ Appendix C MTL Attack Optimization Approximation
As the problem is intractable, we make some approximations and reformulate it to be an Integer Linear Programming (ILP) problem. To do so, we first apply the Taylor expansion on in the numerator of the objective function at the point of :
@@
Why the reminder in Eq. 12 is negligible? is mainly the second-order term in Taylor expansion, which is relatively small compared to the first-order term and thus can be ignored. We provide an analytic upper bound on the second–order remainder and quantify its size relative to the first–order term.
Formally, the second-order term is,
where is the Hessian. Since with the element-wise mask and the budget , we bound the quadratic remainder as,
where .
Then the relative ratio between the second-order and first-order term would be,
Empirically, for the all-shared multi-task models with Deeplab-ResNet34 and over 100 images in the NYUv2 dataset used in Section 4, we measure in Table 5.
| Ratio | Mean Std | Min | Max |
Therefore taking the mean of the empirical results and the common setting of , the ratio , which means the second–order term contributes less than of the linear term, confirming that the linearization error is negligible.
Why we can solve Eq. 11 with LP relaxation? As a commonly-used technique in solving ILP, LP relaxation and rounding do not introduce approximation error for our attack optimization problem [williamson2011design].
Notice that both and in Eq. 11 are pixel-level matrices as the input image , we can decompose it into pixel-level maximization in pixel .
where is the total number of pixels in . Because the objective is linear and separable, the global maximizer is obtained coordinate-wise:
This analytical maximizer matches precisely the algorithm we employ in practice. Starting from an all-zero mask, we iteratively update along the gradient direction , i.e., . This ascent drives every active coordinate to its extreme value, i.e., . The integer solution is then obtained by a simple sign operation, . Accordingly, the LP relaxation followed by sign rounding is exact for DGBA, introducing no discretization error.
Appendix D Auto-SAGA and DGBA-Auto-SAGA
As presented in Section 3 in the main paper, integrating DGBA with any existing attack can be easily accomplished by substituting the single-task gradient, i.e., , with the balanced multi-task counterpart, i.e., . In addition to Algorithm 1 in the main paper, we further illustrate how to integrate DGBA into the multi-model attack algorithm Auto-SAGA.
For Auto-SAGA [rathbun2022game], the original attack formulation is
To integrate DGBA with Auto-SAGA, we normalize the gradient of the objective function and with the objective function value. The updated attack will be,
Appendix E More Experimental Results
This section reports evaluation metrics and more experimental results that are omitted from the main paper.
E.1 Evaluation Metrics
@@
E.2 Attack on Clean Multi-Task Models
@@
Figure 6 illustrates the attack performance on Tiny-Taskonomy with Deeplab-ResNet34. To be consistent with the main paper, we conduct the experiments using different variants of DGBA and the adapted multi-task attacks. The x-axis represents the perturbation bound ranging from 1 to 16, while the y-axis displays the overall Average Relative Performance (ARP, higher-the-better). Overall, DGBA achieves competitive results compared with baselines and always outperforms baselines when the perturbation is small.
@@
We present the full tables of ARP after the attack of all 25 multi-task models trained on NYUv2 with Deeplab-ResNet34 for perturbation bound in Tables 8 and 9. Similarly, Tables 10 and 11 show the attack results for all 15 multi-task models for Tiny-Taskonomy. Overall, DGBA achieves 80% first place (80 out of 100 cases) on NYUv2 and 88.33% (53 out of 60) on Tiny-Taskonomy, demonstrating the effectiveness of adversarial samples from DGBA.